A proton and an electron are fixed in space with a separation of 979 nm.

a. Calculate the electric potential at the midpoint between the two particles.

b. Find the magnitude of the electric field at the same point.

c. What is the direction of the field?
Answer #1

a) V=0 since potential is independent of mass and charge on electron and proton are equal and opposite in magnitude.So V = (kq/r )-( kq/r)=o

b.) E= kq/r^2 by the positive charge and kq/r^2 by the negative charge, so total field will be twice of kq/r^2 and directed towards the negative charge.

that is E=(2*9*10^9*1.6*10^-19)/(489.5*489.5*10^-18)=120195.2 V/m

c) direction towards the electron.